# GraphGDel: Constructing and Learning Graph Representations of Genome-Scale Metabolic Models for Growth-Coupled Gene Deletion Prediction

### v6

@@ Abstract
In genome-scale constraint-based metabolic models, gene deletion strategies are essential for achieving growth-coupled production, where cell growth and target metabolite synthesis occur simultaneously. Despite the inherently networked nature of genome-scale metabolic models, existing computational approaches rely primarily on sequential data and lack graph representations that capture their complex relationships, as both well-defined graph constructions and learning frameworks capable of exploiting them remain largely unexplored. To address this gap, we present a twofold solution. First, we introduce a systematic pipeline for constructing graph representations from constraint-based metabolic models. Second, we develop a deep learning framework that integrates these graph representations with gene and metabolite sequence data to predict growth-coupled gene deletion strategies. Across three metabolic models of varying scale, our approach consistently outperforms established baselines, achieves improvements of 14.04%, 16.26%, and 13.18% in overall accuracy. The source code and example datasets are available at: https://github.com/MetNetComp/GraphGDel.
@@ I INTRODUCTION
Computational approaches are central to metabolic engineering, enabling the systematic design and optimization of microbial strains [otero2021automated, garcia2021metabolic, foster2021building, toya2013flux, pharkya2006optimization]. A key application is computational strain design, which employs mathematical models to simulate cellular metabolism and guide the optimization of target metabolite biosynthesis. Among available modeling frameworks, the constraint-based model is widely used in genome-scale metabolic engineering. It comprises two key components: (1) a metabolic network and (2) gene–protein–reaction (GPR) rules. The metabolic network defines the biochemical transformations among metabolites, catalyzed by enzymes encoded by specific genes. GPR rules, typically expressed as Boolean functions, describe the logical relationships between genes, proteins, and reactions, specifying whether a reaction is active based on gene presence. By altering gene content, the metabolic network can be systematically reprogrammed. One widely applied operation is gene deletion, which disables specific reactions by removing genes that encode essential enzymes. This approach, central to metabolic engineering, enables the rational rewiring of metabolism to enhance the biosynthesis of target compounds.
@@
In constraint-based simulations of gene deletions, the objective is not merely to increase the production of a target metabolite, but to achieve growth-coupled production—a state in which the synthesis of the target metabolite is coupled with cellular growth. This coupling is critical for industrial biotech applications, as microorganism strains with higher growth rates are more likely to dominate during repeated passaging, thereby ensuring stable and sustained metabolite production over time. However, under native metabolic conditions, only a limited number of metabolites naturally exhibit growth-coupling. To extend this property to a broader range of target compounds, it is necessary to design gene deletion strategies that reconfigure the metabolic network to enforce growth-coupled production [vieira2019comparison].
@@
To this end, a range of computational methods have been developed over the past decade to identify gene deletion strategies capable of enabling growth-coupled phenotypes [ranganathan2010optforce, lewis2010omic, yang2011emilio, egen2012truncated, rockwell2013redirector, ohno2014fastpros, gu2016idealknock, tamura2018grid, tamura2023gene]. These strategies have been successfully applied in real-world contexts, facilitating the development of engineered strains for applications, including biofuel production and the industrial biosynthesis of platform chemicals [trinh2008minimal, schneider2020extended, banerjee2020genome]. Despite these advances, calculating gene deletion strategies de novo remains computationally intensive, particularly for genome-scale metabolic models that encompass complex network structures and GPR associations involving numerous genes. In response, recent efforts have focused not only on developing more mathematically efficient algorithms to accelerate these computations, but also on establishing databases to store and manage strain design data, thereby enabling more scalable and effective solutions. One prominent example is MetNetComp database [tamura2023metnetcomp], the first web-based platform to curate gene deletion strategies for growth-coupled production within constraint-based metabolic models. MetNetComp extends the gDel_minRN method [tamura2023gene] to compute gene deletion strategies that enable growth coupling, supporting both minimal and maximal gene deletion configurations. Currently, it hosts a comprehensive repository of over 85,000 gene deletion strategies spanning a wide range of metabolites and multiple constraint-based models from diverse species. Such resources are essential not only for meeting the growing demand for gene deletion analysis but also for enabling new paradigms in strain design, including data-driven approaches to gene deletion strategy development.
@@
Designing data-driven approaches for gene deletion strategy development remains a non-trivial task, with a central challenge being the effective extraction and utilization of information from existing gene deletion strategy databases. The first effort in this direction, DBgDel [yang2025dbgdel], addressed this challenge by summarizing known maximal gene deletion strategies from the database and using them as constraints to limit the search space during optimization. This method strikes a favorable balance between computational efficiency and success rate when applied to genome-scale metabolic models. However, DBgDel relies on hand-crafted extraction rules, limiting its ability to capture complex patterns from large-scale datasets, and still depends on downstream optimization algorithms to properly incorporate GPR rules. To overcome these limitations, DeepGDel [yang2025deepgdel] was introduced as the first framework to formulate gene deletion strategy design as a predictive learning task using deep learning techniques [lecun2015deep], laying the foundation for fully data-driven gene deletion strategy design. DeepGDel jointly models gene and metabolite sequences along with existing gene deletion strategy data, enabling end-to-end prediction of growth-coupled gene deletion strategies and demonstrating robust applicability across metabolic models from diverse species and scales. Although DeepGDel effectively models metabolite features from sequential input data, it ignores upstream and downstream interactions that are widely recognized as essential for understanding the complexity of metabolic systems [structure_1, structure_2, structure_3]. This limitation reduces predictive performance, particularly on large genome-scale models. Graph-based representations provide a natural framework for capturing such structured dependencies and offer critical context for understanding the coupling between target metabolite production and cell growth. Recent advances in graph learning have demonstrated strong potential for extracting meaningful relationships from graph-structured biological data, enabling effective solutions across diverse biological tasks [graphlearning_1, graphlearning_2, graphlearning_3]. Based on these observations, integrating graph-based learning into gene deletion strategy design to better capture metabolite interactions is highly desirable.
@@
However, achieving this goal is not straightforward. Unlike curated pathway maps, such as those provided by KEGG (Kyoto Encyclopedia of Genes and Genomes) [KEGG], genome-scale metabolic models do not explicitly define topological relationships among biological entities in the form of a graph [orth2010flux, simeonidis2015genome]. Instead, they rely on algebraic representations: stoichiometric matrices encoding mass-balance constraints, Boolean expressions defining GPR rules, and flux bounds specified as interval constraints. Since these formulations are inherently algebraic rather than topological, they lack an explicit graph structure in which nodes represent biological entities and edges represent their interactions. This absence of graph structure hinders the direct application of metabolic models in graph-based computation and learning. Moreover, there is a central challenge in constructing meaningful graph representations from metabolic models: the treatment of ubiquitous metabolites, such as ATP, NAD(H), and other common cofactors that participate in a wide range of reactions. These metabolite act as topological hubs, often obscuring the modular organization of metabolic networks and diluting biologically meaningful structures [hubs, structure_3]. Although such entities are commonly collapsed or excluded in databases like KEGG and BiGG to improve visual clarity, their impact on graph-based learning remains underexplored and poorly understood [king2016bigg].
@@
The remaining sections of this paper are organized as follows: Section II-A and Section II-B formalize the fundamental concepts used in this study and illustrate them with a small example. Section III presents (1) the proposed approach for constructing graph representations from constraint-based metabolic models, and (2) the proposed GraphGDel framework together with its implementation details. Section IV-A outlines the setup of the computational experiments. Sections IV-B,IV-C,IV-D, and IV-E describe the datasets, baseline methods, and evaluation metrics. Section IV-F discusses the experimental results, including: (1) performance comparisons between the proposed framework and baseline methods; (2) comparisons of GraphGDel ablation variants that use differernt leraning moudles; and (3) comparisons of GraphGDel variants using different metabolic graph representations across the e_coli_core, iMM904, and iML1515 models. Finally, Section V analyzes the results, evaluates the performance of GraphGDel and its variants, and discusses limitations and directions for future work.
@@ II PRELIMINARY AND PROBLEM SETTING
In this section, we first outline several key terms in this study as a preliminary. Next, we illustrate these concepts using a toy constraint-based model. Finally, we define the main problem in this research, namely, gene deletion strategy prediction. All the notations mentioned in this section are listed in Table I.
TABLE I: Notations used in the definitions in preliminary and problem setting.
Notation Description
$\mathcal{C}$ Constraint-based model
$\mathcal{C}_{1}$ Metabolic network
$\mathcal{C}_{2}$ GPR rule
M Metabolites in the constraint-based model
R Reactions in the constraint-based model
V Reaction rates (flux)
S Stoichiometry matrix
L Lower bounds for reaction rates
U Upper bounds for reaction rates
G Genes in the constraint-based model
F Boolean functions for GPR rule
P Outputs of Boolean functions in F
$m_{\text{target}}$ Target metabolite
$r_{\text{target}}$ Target metabolite production reaction
$r_{\text{growth}}$ Cell growth reaction
$PR_{\text{threshold}}$ Minimum requirement for target production reaction rate
$GR_{\text{threshold}}$ Minimum requirement for cell growth reaction rate
$\textit{G}_{\text{del}}$ Genes deleted in the gene deletion strategy
DS Set of gene deletion strategies
$\textit{DS}_{\text{known}}$ Set of known gene deletion strategies
$\widetilde{\textit{DS}}$ Predicted gene deletion strategies
$\textit{DS}^{*}$ Ground-truth strategies corresponding to $\widetilde{\textit{DS}}$
$\mathcal{F}(\theta)$ Predictive model parameterized by $\theta$
@@ II-A2 Flux balance analysis
When analyzing a metabolic network within a constraint-based model, flux balance analysis (FBA) assumes a steady state in which all metabolic reaction fluxes remain constant [orth2010flux]. Formally, FBA imposes the following conditions: (1) for each metabolite, the sum of producing fluxes equals the sum of consuming fluxes; (2) in each reaction, substrate and product fluxes must satisfy the stoichiometric ratios specified by the reaction equation; and (3) each flux is constrained by prescribed upper and lower bounds. In the standard formulation of FBA applied to a metabolic network $\mathcal{C}_{1}$, the objective is to maximize the cell growth reaction rate $v_{\text{growth}}$, expressed as the following linear program (LP):
@@ II-A3 Growth-coupled production
Growth-coupled production is a special case of microbial metabolism, wherein cell growth co-occurs with the synthesis of a target metabolite. In this study, we adopt the paradigm of weakly growth-coupled production, which requires achieving a non-zero synthesis rate of the target metabolite at the maximum non-zero growth rate [schneider2021systematizing, alter2019determination]. During the simulation of constraint-based models, growth-coupled production can be characterized by two indices: the lower bound of the target metabolite production rate ($PR_{\text{threshold}}$) and the lower bound of the cell growth reaction rate ($GR_{\text{threshold}}$). When the growth rate is maximized, if the simulated values $v_{\text{growth}}$ and minimum $v_{\text{target}}$ exceed their respective lower bounds, growth-coupled production is considered to be achieved. In this study, we set $PR_{\text{threshold}}=GR_{\text{threshold}}=0.001$.
@@ II-B1 Toy constraint-based model setup
Fig. 1 shows a toy example in which $\mathcal{C}=\{\textit{M},\textit{R},\textit{S},\textit{L},\textit{U},\textit{G},\textit{F},\textit{P}\}$ represents a constraint-based model. Specifically, for its metabolic network part $\mathcal{C}_{1}$, we have $\textit{M}=\{m_{1},\ldots,m_{5}\}$, $\textit{R}=\{r_{1},\ldots,r_{8}\}$, $\textit{S}=\begin{pmatrix}1&0&-1&-1&0&0&0&0\\ 0&1&0&0&-1&0&0&0\\ 0&0&1&0&0&-1&0&0\\ 0&0&0&0&1&1&-1&0\\ 0&0&0&1&0&1&0&-1\\ \end{pmatrix}$, $\textit{L}=\{0,0,0,0,0,0,0,0\}$, and $\textit{U}=\{10,10,5,10,10,5,10,10\}$. For its gene reaction rule part $\mathcal{C}_{2}$, we have $\textit{G}=\{g_{1},\ldots,g_{5}\}$, $\textit{F}=\{f_{1},\ldots,f_{8}\}$, and $\textit{P}=\{p_{1},\ldots,p_{8}\}$, where
[…]
@@
TABLE II: Gene deletion strategies for the toy example in Fig. 1 are classified into five types based on the resulting flux (reaction rate) distributions.
@@ 4 $g_{3}$, $g_{4}$}, $g_{3}$, $g_{4}$, $g_{5}$}
TABLE III: The resulting flux distributions for gene deletion strategy Types 1 to 5. For each type, the most optimistic (best) and pessimistic (worst) target metabolite production rate (PR, $v_{8}$) at cell growth rate (GR, $v_{7}$) maximization are provided.
ID
Deletion
strategy
PR
situation
Flux distribution
@@ II-B2 Gene deletions on the toy constraint-based model
Table II describes the patterns of gene deletions in this example: the $2^{5}=32$ patterns are classified into five types based on the flux distributions. The detailed flux distributions for all types can be found in Table III.
@@
If both the path leading to cell growth $(r_{1}\rightarrow r_{5}\rightarrow r_{7})$ and the path leading to target production $(r_{2}\rightarrow r_{4}\rightarrow r_{8})$ are used at maximum flux, then $GR=10$ and $PR=10$ are obtained, as shown in the flux distribution of ID 1 in Table III. This represents the most optimistic case for the flux value of PR. In contrast, if only the path $(r_{1}\rightarrow r_{5}\rightarrow r_{7})$ is used at maximum flux, then $GR=10$ and $PR=0$ are obtained, as shown in the flux distribution of ID 2 in Table III. This represents the most pessimistic case for the flux value of PR. As previously defined, in the context of growth-coupled production, we evaluate PR in the most pessimistic case when GR is maximized. Therefore, in the original state of the toy model, we have $GR=10$ and $PR=0$. We denote this flux distribution as Type 1. Seven gene deletion strategies, $\{g_{1}\}$, $\{g_{2}\}$, $\{g_{3}\}$, $\{g_{1},g_{2}\}$, $\{g_{1},g_{3}\}$, $\{g_{2},g_{3}\}$, and $\{g_{1},g_{2},g_{3}\}$, are also classified as Type 1, as they yield the same flux distribution.
IDs 3 and 4 in Table III describe the most optimistic and pessimistic flux distributions with respect to PR when $g_{4}$ is deleted. Deleting $g_{4}$ gives $f_{4}:p_{4}=g_{4}\wedge g_{5}=0$, forcing $v_{4}=0$. The maximal growth rate is $GR=10$, since all paths to the growth reaction $r_{7}$ (from $r_{1}$ or $r_{2}$ to $r_{7}$) remain available. In the most optimistic case, both the path $r_{1}\rightarrow r_{5}\rightarrow r_{7}$ and the path $r_{2}\rightarrow r_{3}\rightarrow r_{6}\rightarrow r_{7}$ are used, with the latter achieving maximal flux, resulting in $GR=10$ and $PR=5$. In the most pessimistic case, only the first path is used at maximal flux, giving $GR=10$ and $PR=0$. We denote this flux distribution as Type 2. The gene deletion strategies $\{g_{5}\}$, $\{g_{3},g_{5}\}$, and $\{g_{4},g_{5}\}$ are also classified as Type 2, as shown in Table III.
IDs 5 and 6 in Table III describe the flux distributions when $g_{1}$ and $g_{4}$ are deleted. In this case, $f_{3}:p_{3}=g_{1}=0$ and $f_{4}:p_{4}=g_{4}\wedge g_{5}=0$, forcing $v_{3}=0$ and $v_{4}=0$. The maximal growth rate is $GR=10$, since the path $r_{1}\rightarrow r_{5}\rightarrow r_{7}$ remains available. This yields a uniquely determined flux distribution with $v_{1}=v_{5}=10$ and $v_{2}=v_{3}=v_{4}=v_{6}=0$. Consequently, both the optimistic and pessimistic PR values are zero, with $GR=10$. We denote this flux distribution as Type 3. Eleven additional gene deletion strategies are also classified as Type 3, as shown in Table III.
IDs 7 and 8 in Table III describe the flux distributions when $g_{3}$ and $g_{4}$ are deleted. In this case, $f_{4}:p_{4}=g_{4}\wedge g_{5}=0$ and $f_{5}:p_{5}=g_{3}\vee g_{4}=0$, forcing $v_{4}=0$ and $v_{5}=0$. This yields a uniquely determined flux distribution with $v_{2}=v_{3}=v_{6}=4$ and $v_{1}=v_{4}=v_{5}=0$. Consequently, both the optimistic and pessimistic PR values are 5, with maximal $GR=5$. We denote this flux distribution as Type 4. Another gene deletion strategy, $\{g_{3},g_{4},g_{5}\}$, is also classified as Type 4, as shown in Table III.
IDs 9 and 10 in Table III describe the flux distributions when $g_{1}$, $g_{3}$, and $g_{4}$ are deleted. In this case, $f_{3}:p_{3}=g_{1}=0$, $f_{4}:p_{4}=g_{4}\wedge g_{5}=0$, and $f_{5}:p_{5}=g_{3}\vee g_{4}=0$, forcing $v_{3}=v_{4}=v_{5}=0$. The maximal growth rate is $GR=0$, as no path to the growth reaction $r_{7}$ remains viable. This yields a uniquely determined flux distribution with all reaction fluxes equal to zero, resulting in $PR=0$ in both the optimistic and pessimistic cases. We denote this flux distribution as Type 5. The gene deletion strategies $\{g_{2},g_{3},g_{4}\}$, $\{g_{1},g_{2},g_{3},g_{4}\}$, $\{g_{1},g_{3},g_{4},g_{5}\}$, $\{g_{2},g_{3},g_{4},g_{5}\}$, and $\{g_{1},g_{2},g_{3},g_{4},g_{5}\}$ are also classified as Type 5, as shown in Table III.

### v7

@@ Abstract
In genome-scale constraint-based metabolic models, gene deletion strategies are essential for achieving growth-coupled production, where cell growth and target metabolite synthesis occur simultaneously. Despite the inherently networked nature of genome-scale metabolic models, existing computational approaches rely primarily on sequential data and lack graph representations that capture their complex relationships, as both well-defined graph constructions and learning frameworks capable of exploiting them remain largely unexplored. To address this gap, we present a twofold solution. First, we introduce a systematic pipeline for constructing graph representations from constraint-based metabolic models. Second, we develop a deep learning framework that integrates these graph representations with gene and metabolite sequence data to predict growth-coupled gene deletion strategies. Across three metabolic models, our approach consistently outperforms established baselines, with improvements in overall accuracy of 14.04%, 16.26%, and 13.18% over a deep feedforward neural network baseline, 6.17%, 4.96%, and 5.31% over a sequence-learning baseline, and 5.10%, 4.36%, and 4.70% over a topology-aware graph aggregation baseline on the same metabolite graph, respectively. The source code and example datasets are available at: https://github.com/MetNetComp/GraphGDel.
@@ I INTRODUCTION
Computational approaches are central to metabolic engineering, enabling the systematic design and optimization of microbial strains [21, 5, 4, 34, 24]. A key application is computational strain design, which employs mathematical models to simulate cellular metabolism and guide the optimization of target metabolite biosynthesis. Among available modeling frameworks, the constraint-based model is widely used in genome-scale metabolic engineering. It comprises two key components: (1) a metabolic network and (2) gene–protein–reaction (GPR) rules. The metabolic network defines the biochemical transformations among metabolites, catalyzed by enzymes encoded by specific genes. GPR rules, typically expressed as Boolean functions, describe the logical relationships between genes, proteins, and reactions, specifying whether a reaction is active based on gene presence. By altering gene content, the metabolic network can be systematically reprogrammed. One widely applied operation is gene deletion, which disables specific reactions by removing genes that encode essential enzymes. This approach, central to metabolic engineering, enables the rational rewiring of metabolism to enhance the biosynthesis of target compounds.
@@
In constraint-based simulations of gene deletions, the objective is not merely to increase the production of a target metabolite, but to achieve growth-coupled production—a state in which the synthesis of the target metabolite is coupled with cellular growth. This coupling is critical for industrial biotech applications, as microorganism strains with higher growth rates are more likely to dominate during repeated passaging, thereby ensuring stable and sustained metabolite production over time. However, under native metabolic conditions, only a limited number of metabolites naturally exhibit growth-coupling. To extend this property to a broader range of target compounds, it is necessary to design gene deletion strategies that reconfigure the metabolic network to enforce growth-coupled production [37].
@@
To this end, a range of computational methods have been developed over the past decade to identify gene deletion strategies capable of enabling growth-coupled phenotypes [40, 3, 25, 18, 6, 32, 31]. These strategies have been successfully applied in real-world contexts, facilitating the development of engineered strains for applications, including biofuel production and the industrial biosynthesis of platform chemicals [35, 27, 2]. Despite these advances, calculating gene deletion strategies de novo remains computationally intensive, particularly for genome-scale metabolic models that encompass complex network structures and GPR associations involving numerous genes. In response, recent efforts have focused not only on developing more mathematically efficient algorithms to accelerate these computations, but also on establishing databases to store and manage strain design data, thereby enabling more scalable and effective solutions. One prominent example is MetNetComp database [33], the first web-based platform to curate gene deletion strategies for growth-coupled production within constraint-based metabolic models. MetNetComp extends the gDel_minRN method [31] to compute gene deletion strategies that enable growth coupling, supporting both minimal and maximal gene deletion configurations. Currently, it hosts a comprehensive repository of over 85,000 gene deletion strategies spanning a wide range of metabolites and multiple constraint-based models from diverse species. Such resources are essential not only for meeting the growing demand for gene deletion analysis but also for enabling new paradigms in strain design, including data-driven approaches to gene deletion strategy development.
@@
Designing data-driven approaches for gene deletion strategy development remains a non-trivial task, with a central challenge being the effective extraction and utilization of information from existing gene deletion strategy databases. The first effort in this direction, DBgDel [41], addressed this challenge by summarizing known maximal gene deletion strategies from the database and using them as constraints to limit the search space during optimization. This method strikes a favorable balance between computational efficiency and success rate when applied to genome-scale metabolic models. However, DBgDel relies on hand-crafted extraction rules, limiting its ability to capture complex patterns from large-scale datasets, and still depends on downstream optimization algorithms to properly incorporate GPR rules. To overcome these limitations, DeepGDel [42] was introduced as the first framework to formulate gene deletion strategy design as a predictive learning task using deep learning techniques [12], laying the foundation for fully data-driven gene deletion strategy design. DeepGDel jointly models gene and metabolite sequences along with existing gene deletion strategy data, enabling end-to-end prediction of growth-coupled gene deletion strategies and demonstrating robust applicability across metabolic models from diverse species and scales. Although DeepGDel effectively models metabolite features from sequential input data, it ignores upstream and downstream interactions that are widely recognized as essential for understanding the complexity of metabolic systems [29, 38, 8]. This limitation reduces predictive performance, particularly on large genome-scale models. Graph-based representations provide a natural framework for capturing such structured dependencies and offer critical context for understanding the coupling between target metabolite production and cell growth. Recent advances in graph learning have demonstrated strong potential for extracting meaningful relationships from graph-structured biological data, enabling effective solutions across diverse biological tasks [13, 17, 9]. Based on these observations, integrating graph-based learning into gene deletion strategy design to better capture metabolite interactions is highly desirable.
@@
However, achieving this goal is not straightforward. Unlike curated pathway maps, such as those provided by KEGG (Kyoto Encyclopedia of Genes and Genomes) [10], genome-scale metabolic models do not explicitly define topological relationships among biological entities in the form of a graph [20, 28]. Instead, they rely on algebraic representations: stoichiometric matrices encoding mass-balance constraints, Boolean expressions defining GPR rules, and flux bounds specified as interval constraints. Since these formulations are inherently algebraic rather than topological, they lack an explicit graph structure in which nodes represent biological entities and edges represent their interactions. This absence of graph structure hinders the direct application of metabolic models in graph-based computation and learning. Moreover, there is a central challenge in constructing meaningful graph representations from metabolic models: the treatment of ubiquitous metabolites, such as ATP, NAD(H), and other common cofactors that participate in a wide range of reactions. These metabolite act as topological hubs, often obscuring the modular organization of metabolic networks and diluting biologically meaningful structures [23, 8]. Although such entities are commonly collapsed or excluded in databases like KEGG and BiGG to improve visual clarity, their impact on graph-based learning remains underexplored and poorly understood [11].
@@ II PRELIMINARY AND PROBLEM SETTING
In this section, we first outline several key terms in this study as a preliminary. Next, we illustrate these concepts using a toy constraint-based model. Finally, we define the main problem in this research, namely, gene deletion strategy prediction. All the notations mentioned in this section are listed in a table, which can be found in the Appendix.
@@ II-A2 Flux balance analysis
When analyzing a metabolic network within a constraint-based model, flux balance analysis (FBA) assumes a steady state in which all metabolic reaction fluxes remain constant [20]. Formally, FBA imposes the following conditions: (1) for each metabolite, the sum of producing fluxes equals the sum of consuming fluxes; (2) in each reaction, substrate and product fluxes must satisfy the stoichiometric ratios specified by the reaction equation; and (3) each flux is constrained by prescribed upper and lower bounds. In the standard formulation of FBA applied to a metabolic network $\mathcal{C}_{1}$, the objective is to maximize the cell growth reaction rate $v_{\text{growth}}$, expressed as the following linear program (LP):
@@ II-A3 Growth-coupled production
Growth-coupled production is a special case of microbial metabolism, wherein cell growth co-occurs with the synthesis of a target metabolite. In this study, we adopt the paradigm of weakly growth-coupled production, which requires achieving a non-zero synthesis rate of the target metabolite at the maximum non-zero growth rate [26, 1]. During the simulation of constraint-based models, growth-coupled production can be characterized by two indices: the lower bound of the target metabolite production rate ($PR_{\text{threshold}}$) and the lower bound of the cell growth reaction rate ($GR_{\text{threshold}}$). When the growth rate is maximized, if the simulated values $v_{\text{growth}}$ and minimum $v_{\text{target}}$ exceed their respective lower bounds, growth-coupled production is considered to be achieved. In this study, we set $PR_{\text{threshold}}=GR_{\text{threshold}}=0.001$.
@@ II-B1 Toy constraint-based model setup
Fig. 1 shows a toy example in which $\mathcal{C}=\{\textit{M},\textit{R},\textit{S},\textit{L},\textit{U},\textit{G},\textit{F},\textit{P}\}$ represents a constraint-based model. Specifically, for its metabolic network part $\mathcal{C}_{1}$, we have $\textit{M}=\{m_{1},\ldots,m_{5}\}$, $\textit{R}=\{r_{1},\ldots,r_{8}\}$, $\textit{S}=\begin{pmatrix}1&0&-1&-1&0&0&0&0\\ 0&1&0&0&-1&0&0&0\\ 0&0&1&0&0&-1&0&0\\ 0&0&0&0&1&1&-1&0\\ 0&0&0&1&0&1&0&-1\\ \end{pmatrix}$, $\textit{L}=\{0,0,0,0,0,0,0,0\}$, and $\textit{U}=\{10,10,5,10,10,5,10,10\}$. For its gene reaction rule part $\mathcal{C}_{2}$, we have $G=\{g_{1},\ldots,g_{5}\}$, $F=\{f_{1},\ldots,f_{8}\}$ and $P=\{p_{1},\ldots,p_{8}\}$, where $f_{i}:p_{i}=1$ for $i\in\{1,2,7,8\}$; $f_{3}:p_{3}=g_{1}$; $f_{4}:p_{4}=g_{4}\wedge g_{5}$; $f_{5}:p_{5}=g_{3}\vee g_{4}$; and $f_{6}:p_{6}=g_{2}$.
@@
TABLE I: Gene deletion strategies for the toy example in Fig. 1 are classified into five types based on the resulting flux (reaction rate) distributions.
@@ 4 $g_{3}$, $g_{4}$}, $g_{3}$, $g_{4}$, $g_{5}$}
TABLE II: The resulting flux distributions for gene deletion strategy Types 1 to 5. For each type, the most optimistic (best) and pessimistic (worst) target metabolite production rate (PR, $v_{8}$) at cell growth rate (GR, $v_{7}$) maximization are provided.
ID Deletion strategy PR situation Flux distribution
@@ II-B2 Gene deletions on the toy constraint-based model
Table I describes the patterns of gene deletions in this example: the $2^{5}=32$ patterns are classified into five types based on the flux distributions. The detailed flux distributions for all types can be found in Table II.
@@
If both the path leading to cell growth $(r_{1}\rightarrow r_{5}\rightarrow r_{7})$ and the path leading to target production $(r_{2}\rightarrow r_{4}\rightarrow r_{8})$ are used at maximum flux, then $GR=10$ and $PR=10$ are obtained, as shown in the flux distribution of ID 1 in Table II. This represents the most optimistic case for the flux value of PR. In contrast, if only the path $(r_{1}\rightarrow r_{5}\rightarrow r_{7})$ is used at maximum flux, then $GR=10$ and $PR=0$ are obtained, as shown in the flux distribution of ID 2 in Table II. This represents the most pessimistic case for the flux value of PR. As previously defined, in the context of growth-coupled production, we evaluate PR in the most pessimistic case when GR is maximized. Therefore, in the original state of the toy model, we have $GR=10$ and $PR=0$. We denote this flux distribution as Type 1. Seven gene deletion strategies, $\{g_{1}\}$, $\{g_{2}\}$, $\{g_{3}\}$, $\{g_{1},g_{2}\}$, $\{g_{1},g_{3}\}$, $\{g_{2},g_{3}\}$, and $\{g_{1},g_{2},g_{3}\}$, are also classified as Type 1, as they yield the same flux distribution.
IDs 7 and 8 in Table II describe the flux distributions when $g_{3}$ and $g_{4}$ are deleted. In this case, $f_{4}:p_{4}=g_{4}\wedge g_{5}=0$ and $f_{5}:p_{5}=g_{3}\vee g_{4}=0$, forcing $v_{4}=0$ and $v_{5}=0$. This yields a uniquely determined flux distribution with $v_{2}=v_{3}=v_{6}=4$ and $v_{1}=v_{4}=v_{5}=0$. Consequently, both the optimistic and pessimistic PR values are 5, with maximal $GR=5$. We denote this flux distribution as Type 4. Another gene deletion strategy, $\{g_{3},g_{4},g_{5}\}$, is also classified as Type 4, as shown in Table II.
The detailed analysis of the flux distribution under deletion strategies Type 2, 3, and 5 can be found in the Appendix.